OpenStudy (anonymous):

show that sec^2 x/2 = 2/ 1+cos x

OpenStudy (anonymous):

\[\cos2\theta=cos^2\theta-sin^2\theta\] Use this..

OpenStudy (anonymous):

ok i will try.thanks anyway :)

OpenStudy (anonymous):

the last part i got stuck in 4/ 1+cos 2x.

OpenStudy (anonymous):

\[cos2\theta = 2cos^2\theta - 1 \] \[cos2\theta + 1 = 2cos^2\theta\] \[\frac{1}{sec\theta}=cos\theta\]

OpenStudy (anonymous):

Now you'll get it for sure (y) :)...

OpenStudy (anonymous):

yeah i got it :) thanks dude.